OpenStudy (anonymous):

Solve for x in the proportion x over the quantity x plus 12 equals 1 over the quantity x minus 1. x ≈ −6.02 and x ≈ 5.02 x ≈ −3.46 and x ≈ 3.46 x ≈ −3.14 and x ≈ 4.14 x ≈ −2.61 and x ≈ 4.61

OpenStudy (anonymous):

x/(x+12)=1/(x-1) x(x-1)=x+12 x^2-x=x+12 x^2-x-x-12=0 x^2-2x-12=0

OpenStudy (anonymous):

x/(x+12)=1/(x-1) x(x-1)=x+12 x^2-x=x+12 x^2-x-x-12=0 x^2-2x-12=0